OpenStudy (anonymous):

Help/Equations A savings account was opened with an initial deposit and never touched again. An account that does not meet the minimum balance requirement is charged a service fee. Assuming the account was open with an initial deposit below the minimum amount, the current account balance, B, is equal to the product of the initial deposit, I, and the quantity of the difference of 1 and the monthly service fee, f, raised to the power of n, the number of months since the account was open. If an initial deposit of $2,228.00 is made to the account with a monthly service fee of 1%, what will be the value of B after 21 months? Round to the nearest cent. $2,745.77 $1,760.12 $1,804.07 $1,822.30

OpenStudy (anonymous):

B = I ( 1 - f^n) This is what the formula they were referring to I believe

OpenStudy (anonymous):

Given I = $ 2,228.00 f = 1% which is 1/100 n = 21 months

OpenStudy (anonymous):

thanks @ganibl for helping me seems like the people i thought would help wont so thanks for sticking with me through these equations

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

so what am i suppose to do after that

OpenStudy (anonymous):

Just Substitute those values in the formula and simplify it u will get the answer

OpenStudy (anonymous):

so wouldnt it make it C

OpenStudy (anonymous):

I am getting same answer as $ 2,228. Not sure if I misunderstood the problem or some where I might have gone wrong

OpenStudy (anonymous):

Ha got it

OpenStudy (anonymous):

B = I * ( 1 -f )^n B = $2,228 * ( 1 - 1/100)^21 B =$2228 * ( 0.99)^21 = $1804.07 Cents So U r Right

OpenStudy (anonymous):

Sorry I took the wrong formula in the beginning